OpenStudy (anonymous):

factor 15x^3-6x^2-25x+10 by grouping

OpenStudy (anonymous):

Factor each 3x^2(5x-2)-5(5x-2)=(3x^2-5)(5x-2). If you need to take it further, (3x^2-5) factors into \[3(x^2-5/3)=3(x-\sqrt(5/3))(x+\sqrt(5/3))\]

OpenStudy (gw2011):

(15x^3-6x^2) + (-25x+10) 3x^2(5x-2)-5(5x-2) (3x^2-5)(5x-2)